# HiT: History-Injection Transformers for Onboard Continuous Flood Change Detection

### v2

@@ 6.1 Memory footprint
6.2 Temporal persistance
Unfortunately, we do not have a test dataset for continuous change detection for natural disasters. Therefore, we cannot directly compute how much information learned from previous time steps is used in the inference process. Notably, the bitemporal baseline model can be used to identify low-quality images in the test series, in which case we expect our model to utilize information stored in the HE. We found candidates in the following way. Each pre-disaster image in the input series is paired with the post-disaster image for inference by the bitemporal baseline. If the baseline performance on a pair of images is much worse than on the other pairs, then the particular pre-disaster image must be of a bad quality. By observing the difference in F1-score between the best result and each pre-disaster image, we identified two primary candidates for low-quality images: Brazil No. 3 and Germany No. 3 . The complete analysis is shown in Table 4 in Appendix 0.A and images of interest can be seen in Appendix 0.B in Fig. 9 and Fig. 8.
Our HiT-Prithvi model was evaluated analogously. During series processing, it was prompted to create a prediction for the post-disaster image. As shown in Table 3, the HiT model did not experience any drop in performance when presented with low-quality images before the final prediction. This suggests its capability to store important information in the HE. However, this requires further investigation, which is not possible without specialized testing data.
@@ 7 Conclusion
This work addresses the challenge of continuous onboard change detection for natural disaster monitoring by introducing the History Injection mechanism for Transformers (HiT). Our proposed HiT-Prithvi model integrates compact History Embeddings directly into the Prithvi-EO-2.0-tiny encoder, enabling efficient multi-temporal flood detection under the strict constraints of small satellite platforms. The HiT mechanism reduces storage requirements by 99.61% compared to storing raw imagery, allowing continental-scale monitoring with only a few gigabytes of memory. Despite this dramatic compression, the model achieves an F1-score of 0.479 on the STTORM-CD test set—comparable to bi-temporal baselines—while producing almost no false positives on unchanged scenes. Real-time inference at 43 FPS on Jetson Orin Nano demonstrates the model’s practical viability for onboard deployment. Our experiments reveal that the model is robust to variations in fusion stage, embedding dimensionality, and spatial resolution, with optimal performance achieved using 64 tokens of dimension 24 and fusion stage 5. Preliminary results suggest that the History Embedding successfully retains temporal information, maintaining performance even when intermediate observations are degraded.
@@ Appendix 0.A Baseline low-quality image identification
Table 4 shows the mean standard deviation of the difference from the best result in the series.
Table 4: Baseline F1 difference to the best result.
@@
Appendix 0.B Models Test Results
This section contains segmentation outputs for the entire STTORM-CD test set for both HiT-Prithvi and the bitemporal baseline. Segmentation results are shown as RGB images, with green for true positive, red for false positive, blue for false negative, and black for true negative. The test set consists of 4 scenes: Germany in Fig. 8, Brazil in Fig. 9, Laos in Fig. 7, and Niger in Fig. 10. In the Niger scene, both models perform poorly. Not a single instance of the bitemporal baseline was able to detect any change in this scene. In contrast, the HiT model reports some change in this case, and noticeable activity is visible in the raw probabilities.
Figure 8: Model’s results for the Germany test scene. Top line: the input series images. Second and Third row: segmentation output of HiT-Prithvi model in raw probabilities and probabilities >0.5>0.5, respectively. The last row: segmentation output of bitemporal baseline in probabilities >0.5>0.5. (Green = TP, Red = FP, Blue = FN, Black = TN).
Figure 9: Model’s results for the Brazil test scene. Top line: the input series images. Second and Third row: segmentation output of HiT-Prithvi model in raw probabilities and probabilities >0.5>0.5, respectively. The last row: segmentation output of bitemporal baseline in probabilities >0.5>0.5. (Green = TP, Red = FP, Blue = FN, Black = TN).
Figure 10: Model’s results for the Niger test scene. Top line: the input series images. Second and Third row: segmentation output of HiT-Prithvi model in raw probabilities and probabilities >0.5>0.5, respectively. The last row: segmentation output of bitemporal baseline in probabilities >0.5>0.5. (Green = TP, Red = FP, Blue = FN, Black = TN).

### v3

@@ 6.1 Resource Efficiency: Memory footprint
6.2 Temporal Persistence and Prediction Stability
To evaluate the model’s capacity to continuously store historical information, we designed an experiment to observe how it leverages information from the HE during the processing of low-quality inputs. To isolate these low-quality images, we employed a bi-temporal baseline model. In this setup, each pre-disaster image in a series was individually paired with the post-disaster image for inference. The underlying logic is straightforward: if the baseline model yields a substantially lower performance on a specific pair compared to the rest of the series, that particular pre-disaster image is deemed to be of low quality. By analyzing the drop in F1-score between the best overall result and each individual pre-disaster image, we successfully identified two primary low-quality candidates: Brazil No. 3 and Germany No. 3. The comprehensive analysis can be found in Table 5 of Appendix 0.A, with the corresponding test scenes depicted in Fig. 8.
Figure 8: Test scenes from STTORM-CD-Floods [herec2025sttorm] (a) Brazil, (b) Germany.
@@ 6.2 Temporal Persistence and Prediction Stability
Notably, while Brazil No. 1 produced the worst score, it could not be utilized for this test. Because it is the first image in the series, the HE is empty at that moment, precluding any measurement of continuous memory. Our HiT-PrithVi model was evaluated analogously. During series processing, it was prompted to create a prediction for the post-disaster image. As shown in Table 4, the HiT model did not experience any drop in performance when presented with low-quality images before the final prediction. This suggests its capability to store important information in the HE. However, this requires further investigation, which is not possible without specialized testing data.
6.3 Error Analysis
The model failure can be seen on the Niger test scene from STTORM-CD-Floods in Fig. 9. This scenario tests the model‘s generalization on biome not present in the training set. HiT-Pritvi was able to identify only a small region with change, while the baseline model failed completely. The reference ContUrbanCD model detects change precisely with the cost of false positive detections in the timesteps with no change. This problem is even more prevalent on the Brazil scene in Fig. 10. Although both baseline and HiT-PrithVi have correctly predicted no change in first 3 predictions, ContUrbanCD has again produced false positive change predictions. For real life use cases, HiT-PrithVi behavior is preferable to ContUrbanCD, as false positive detections in no change image will trigger costly downlink of the scene.
Figure 9: HiT-PrithVi, bi-temporal baseline, ContUrbanCD results for the Niger scene.Segmentation results: TP (green), FP (red), TN (blue), TN (black).
Figure 10: HiT-PrithVi, bi-temporal baseline, ContUrbanCD results for the Brazil scene. Segmentation results: TP (green), FP (red), TN (blue), TN (black).
@@ 7 Conclusion
This work addresses the challenge of continuous onboard change detection for natural disaster monitoring by introducing the History Injection mechanism for Transformers (HiT). Our proposed HiT-PrithVi model integrates compact History Embeddings directly into the PrithVi-EO-2.0-tiny encoder, enabling efficient multi-temporal flood detection under the strict constraints of small satellite platforms. The HiT mechanism reduces storage requirements by 99.61% compared to storing raw imagery, allowing continental-scale monitoring with only a few gigabytes of memory. Despite this dramatic compression, the model achieves an F1-score of 0.479 on the STTORM-CD-Floods test set—comparable to bi-temporal baselines—while producing almost no false positives on scenes without change. Real-time inference at 43 FPS on Jetson Orin Nano 8GB demonstrates the model’s practical viability for onboard deployment. Our experiments reveal that the model is robust to variations in fusion stage, embedding dimensionality, and spatial resolution, with optimal performance achieved using 64 tokens of dimension 24 and injection depth 5. Analysis on the STTORM-CD-Floods Brazil scene suggests that the History Embedding successfully retains temporal information, maintaining performance even when intermediate observations are degraded.
@@ Appendix 0.A Baseline low-quality image identification
Table 5 shows the mean standard deviation of the difference from the best result in the series for bi-temporal baseline.
Table 5: Bi-temporal baseline F1 difference to the best result.